If M and N are negative integers and -3M+4N=10,Which of the following could be the value of N?
a. -2
b. -4
c.-6
d. -7
e. -9

Answer by JoelSchwartz(130)

If N=-2
-3M+4*-2=10
-3M-8=10
-3M=18
-M=6
M=-6
The only N value listed that gives a integer for M is -2.